# dot-PE: Sampler-free gravitational wave inference using matrix multiplication

### v1

@@ 5.2 Parameter Estimation
To evaluate the performance and bias of parameter estimation runs, we use dot-PE to sample the posteriors of the synthetic injections describe in the beginning of this chapter. We impose the same $\langle h\mid h\rangle\in(70,200)$ criterion on the posterior samples used to generate the plots. The parameter estimation runs are performed with intrinsic samples banks of $N_{\rm int.}=2^{16}$, $N_{\rm ext.}=2^{10}$ extrinsic samples and $N_{\phi}=64$. A few percent of the events resulted in less than 30 samples (overall, or in the allowed $\langle h\mid h\rangle$ range) and are excluded from the analysis. Per injection and set of samples, we evaluate the cumulative density functions (CDF) of several well-measured parameters. The evaluated CDF values are expected to distribute uniformly in $(0,1)$. The resulting probability–probability (P–P) plots are shown in Fig. 6, where agreement between the empirical CDFs (solid colored lines) and the ideal diagonal (black dotted lines) indicates the accuracy of the method.
@@ 6 Conclusions
We have presented a novel approach to the posterior estimation problem of CBC GW events. We split the 15 dimensions of parameter space into 4 subsets, according to their functional form in the likelihood function. This allows us to evaluate the (distance-marginalized) likelihood for each intrinsic-extrinsic-phase combination, using a sequences of matrix multiplications. We implemented this method in dot-PE, a pure Python package. The method allows us to integrate the Bayesian evidence ratio, to accuracy of ±$\pm 1$ in $\ln\mathcal{Z}$ within a few minutes (on a single CPU core). The runtime can be greatly reduced by trivial parallelization (already implemented in the current code) and GPU usage. The relations between the number of samples used, $N_{\rm int.}$, $N_{\rm ext.}$ and $N_{\phi}$ to the runtime and performance of the integration or parameter estimation is predictable and well understood, as demonstrated in Section 5 and Appendix D. For the purpose of a search, dot-PE could produce the evidence ratio, which is the optimal test statistic and includes the effects of precession and higher modes, for many thousands of candidates provided by a traditional template–bank search.
@@ Appendix C Extrinsic parameter marginalization
Where $\beta\sim 0.5$ is an inverse-temperature regularization term, pre-optimized to maximize the effective sample size. A large number $N_{\mathrm{QMC}}$ of arrival-time samples $\left\{\bm{\tau}_{q}\right\}_{q=1}^{N_{\mathrm{QMC}}}$ are drawn from the distributions (through their cumulatives). The exclusion of non-physical arrival times inconsistent with any sky position is discussed in detail in [25], but omitted here. The samples of arrival time at each detector are then translated to a global arrival times $t_{\oplus}$ and sky-positions $\lambda,\varphi$, by drawing from all sky-position samples consistent (within the time resolution specified) with the inter-detector time delays. Since the intrinsic components of the waveforms are already known, it is relatively quick to evaluate the likelihood components (adopting the notations used in Section 4.2) $\langle d\mid h\rangle_{em}$ and $\langle h\mid h\rangle_{em}$ (for the given intrinsic samples and all extrinsic samples and modes), then apply different $\phi_{\rm ref}$, marginalize over distance, and average over phase for ¯$\overline{\mathcal{L}}_{\phi d_{L}}(\bm{\theta}_{\rm ext}|\bm{\theta}_{\rm int})$. The marginalized likelihood of the intrinsic sample can now be evaluated as
¯$\overline{\mathcal{L}}(\bm{\theta}_{\rm int})=\frac{1}{N_{\mathrm{QMC}}}\sum_{e=1}^{N_{\mathrm{QMC}}}\frac{\pi\left(\bm{\theta}_{\rm ext}^{(e)}\right)}{\pi^{\prime}\left(\bm{\theta}_{\rm ext}^{(e)}\right)}\overline{\mathcal{L}}_{\phi d_{L}}(\bm{\theta}_{\rm ext}^{(e)}|\bm{\theta}_{\rm int})$ (72)
@@ Appendix D Complexity and runtime
To test the runtime of the code in realistic settings, we chose $N_{\rm int.}$ and $N_{\rm ext.}$ per intrinsic sample bank based on the results of Section 5.1. In each of $N_{\rm int.}$ and $N_{\rm ext.}$, we selected the minimal value so that the median $\ln\mathcal{Z}$ accuracy will be ±$\pm 1$. A single run was performed on the 1,024 injections per mass range described in Section 5. The runs were performed on a single CPU core from the Xeon Gold 6226R CPU with clock rate of 2.10 GHz. As in Section 5.1, the errors are measured relative to the $\ln\mathcal{Z}$ computed using separate denser banks of size $N_{\rm int.}=2^{18}$, and with $N_{\rm ext.}=2^{10}$. The runtimes and typical dimensions are shown in Table 1. We also analyze the expected computational cost (in floating point operations, FLOPs) in Table 2. We do not include cogwheel implemented operations, namely drawing extrinsic samples and evaluating the distance-marginalized likelihood. All operations not appearing in Table 2 are regarded as “Overhead” in Table 1. The core dot-PE linear-algebra operations are dominated in runtime and FLOP count by the pre-selection process. The fractional costs of the pre-selection and overhead appear in Table 1, and they suggest the lower limits of runtime under improvements to either the overhead or to the linear-algebra functions (e.g., using GPUs or a large number of CPUs.).
@@
$N_{\rm int.}^{\prime}/N_{\rm int.}$ $0.01^{0.03}_{0.005}$ $0.02^{0.05}_{0.01}$ $0.1^{0.1}_{0.02}$ $0.2^{0.3}_{0.1}$ $0.6^{0.8}_{0.4}$ $0.3^{0.4}_{0.2}$
$N_{\rm margin.}/\left(N_{\rm int.}^{\prime}N_{\rm ext.}N_{\phi}\right)$ $0.01^{0.01}_{0.003}$ $0.01^{0.01}_{0.005}$ $0.02^{0.03}_{0.01}$ $0.03^{0.1}_{0.02}$ $0.1^{0.1}_{0.03}$ $0.2^{0.7}_{0.1}$
$\left|\Delta\ln\mathcal{Z}\right|$ $0.52^{1.07}_{0.20}$ $0.45^{0.93}_{0.21}$ $0.58^{1.12}_{0.28}$ $0.59^{1.10}_{0.29}$ $0.55^{1.07}_{0.23}$ $0.68^{1.22}_{0.31}$
@@
Pre-selection/Total runtime $0.81^{0.82}_{0.79}$ $0.81^{0.83}_{0.79}$ $0.54^{0.56}_{0.52}$ $0.14^{0.14}_{0.13}$ $0.04^{0.04}_{0.04}$ $0.13^{0.14}_{0.12}$
@@
Table 1: Summary of runs aimed $\ln\mathcal{Z}$ accuracy of ±$\pm 1$. $N_{\rm int.}$ and $N_{\rm ext.}$ are chosen based on the accuracies presented in Fig. 2. Subscripts and sup-scripts stand for 0.25 and 0.75 quantile, based on 1,024 injections. Runtimes quoted correspond to using a single CPU, see body of text.

### v3

@@ 5.2 Parameter Estimation
To evaluate the performance and bias of parameter estimation runs, we use dot-PE to sample the posteriors of the synthetic injections describe in the beginning of this section. We impose the same $\langle h\mid h\rangle\in(70,200)$ criterion on the posterior samples used to generate the plots. The parameter estimation runs are performed with intrinsic samples banks of $N_{\rm int.}=2^{16}$, $N_{\rm ext.}=2^{10}$ extrinsic samples and $N_{\phi}=64$. A few percent of the events resulted in less than 30 samples (overall, or in the allowed $\langle h\mid h\rangle$ range) and are excluded from the analysis. Per injection and set of samples, we evaluate the cumulative density functions (CDF) of several well-measured parameters. The evaluated CDF values are expected to distribute uniformly in $(0,1)$. The resulting probability–probability (P–P) plots are shown in Fig. 6, where agreement between the empirical CDFs (solid colored lines) and the ideal diagonal (black dotted lines) indicates the accuracy of the method.
@@ 6 Conclusions
We have presented a novel approach to the posterior estimation problem of CBC GW events. We split the 15 dimensions of parameter space into 4 subsets, according to their functional form in the likelihood function. This allows us to evaluate the (distance-marginalized) likelihood for each intrinsic-extrinsic-phase combination, using a sequences of matrix multiplications. We implemented this method in dot-PE, a pure Python package. The method allows us to integrate the Bayesian evidence ratio, to accuracy of $\pm 1$ in $\ln\mathcal{Z}$ within a few minutes (on a single CPU core). The runtime can be greatly reduced by trivial parallelization (already implemented in the current code) and GPU usage. The relations between the number of samples used, $N_{\rm int.}$, $N_{\rm ext.}$ and $N_{\phi}$ to the runtime and performance of the integration or parameter estimation is predictable and well understood, as demonstrated in Section 5 and Appendix D. For the purpose of a search, dot-PE could produce the evidence ratio, which is the optimal test statistic and includes the effects of precession and higher modes, for many thousands of candidates provided by a traditional match-filtering search.
@@ Appendix C Extrinsic parameter marginalization
Where $\beta\sim 0.5$ is an inverse-temperature regularization term, pre-optimized to maximize the effective sample size. A large number $N_{\mathrm{QMC}}$ of arrival-time samples $\left\{\bm{\tau}_{q}\right\}_{q=1}^{N_{\mathrm{QMC}}}$ are drawn from the distributions (through their cumulatives). The exclusion of non-physical arrival times inconsistent with any sky position is discussed in detail in [25], but omitted here. The samples of arrival time at each detector are then translated to a global arrival times $t_{\oplus}$ and sky-positions $\lambda,\varphi$, by drawing from all sky-position samples consistent (within the time resolution specified) with the inter-detector time delays. Since the intrinsic components of the waveforms are already known, it is relatively quick to evaluate the likelihood components (adopting the notations used in Section 4.2) $\langle d\mid h\rangle_{em}$ and $\langle h\mid h\rangle_{em}$ (for the given intrinsic samples and all extrinsic samples and modes), then apply different $\phi_{\rm ref}$, marginalize over distance, and average over phase for $\overline{\mathcal{L}}_{\phi d_{L}}(\bm{\theta}_{\rm ext}|\bm{\theta}_{\rm int})$. The marginalized likelihood of the intrinsic sample can now be evaluated as
$\overline{\mathcal{L}}(\bm{\theta}_{\rm int})=\frac{1}{N_{\mathrm{QMC}}}\sum_{e=1}^{N_{\mathrm{QMC}}}\frac{\pi\left(\bm{\theta}_{\rm ext}^{(e)}\right)}{\pi^{\prime}\left(\bm{\theta}_{\rm ext}^{(e)}\right)}\overline{\mathcal{L}}_{\phi d_{L}}(\bm{\theta}_{\rm ext}^{(e)}|\bm{\theta}_{\rm int})$ (72)
@@ Appendix D Complexity and runtime
To test the runtime of the code in realistic settings, we chose $N_{\rm int.}$ and $N_{\rm ext.}$ per intrinsic sample bank based on the results of Section 5.1. In each of $N_{\rm int.}$ and $N_{\rm ext.}$, we selected the minimal value so that the median $\ln\mathcal{Z}$ accuracy will be $\pm 1$. A single run was performed on the 1,024 injections per mass range described in Section 5. The runs were performed on a single CPU core from the Xeon Gold 6226R CPU with clock rate of 2.10 GHz. As in Section 5.1, the errors are measured relative to the $\ln\mathcal{Z}$ computed using separate denser banks of size $N_{\rm int.}=2^{18}$, and with $N_{\rm ext.}=2^{10}$. The runtimes and typical dimensions are shown in Table 1. We also analyze the expected computational cost (in floating point operations, FLOPs) in Table 2. We do not include cogwheel implemented operations, namely drawing extrinsic samples and evaluating the distance-marginalized likelihood. All operations not appearing in Table 2 are regarded as “Overhead” in Table 1. The core dot-PE linear-algebra operations are dominated in runtime and FLOP count by the pre-selection process. The fractional costs of the pre-selection and overhead appear in Table 1, and they suggest the lower limits of runtime under improvements to either the overhead or to the linear-algebra functions (e.g., using GPUs or a large number of CPUs.).
@@
$N_{\rm int.}^{\prime}/N_{\rm int.}$ $01^{0.03}_{0.005} 0.0$$02^{0.05}_{0.01}$ $0.1^{0.1}_{0.02}$ $0.2^{0.3}_{0.1}$ $0.6^{0.8}_{0.4}$ $0.3^{0.4}_{0.2}$
$N_{\rm margin.}/\left(N_{\rm int.}^{\prime}N_{\rm ext.}N_{\phi}\right)$ $01^{0.01}_{0.003} 0.0$$1^{0.01}_{0.005} 0.0$$2^{0.03}_{0.01} 0.0$$03^{0.1}_{0.02}$ $0.1^{0.1}_{0.03}$ $0.2^{0.7}_{0.1}$
$\left|\Delta\ln\mathcal{Z}\right|$ $0.52^{1.07}_{0.20}$ $0.45^{0.93}_{0.21}$ $0.58^{1.12}_{0.28}$ $9^{1.10}_{0.29} 0.550$$5^{1.07}_{0.23}$ $0.68^{1.22}_{0.31}$
@@
Pre-selection/Total runtime $0.81^{0.82}_{0.79}$ $0.81^{0.83}_{0.79}$ $0.54^{0.56}_{0.52}$ $14^{0.14}_{0.13} 0.0$$04^{0.04}_{0.04}$ $0.13^{0.14}_{0.12}$
@@
Table 1: Summary of runs aimed $\ln\mathcal{Z}$ accuracy of $\pm 1$. $N_{\rm int.}$ and $N_{\rm ext.}$ are chosen based on the accuracies presented in Fig. 2. Subscripts and sup-scripts stand for 0.25 and 0.75 quantile, based on 1,024 injections. Runtimes quoted correspond to using a single CPU, see body of text.